Problem 1 (10 marks): Find the Thevenin equivalent at terminals a-b of the following circuit: +- 10 kΩ 20 kΩ 70ν (Ξ)
Answer #1

Denote redraw the loop current and the venin voltage the circuit as shown in Figure 1. + - I /ь 2012 log.n + 4V 7ov ☺ from Fio VE 70 - (10x103) (1x103) = 70-10 Hence, the value of Thevenin bov] voltage, von is since the circuit has 4V depen dent sourcalculate the value of current is ie -V. 10x103 substitute -Iv for V 10x103 loop mnopqr in e 0.1mA apply kirchhoffs voltagesubstitute 0.35 mA for ia in equation (2) Rin = 35x 103 = 2857 · 143 12 = 2.86 Hence, the value of the venin resistance, Riho